Child's Play Company makes a plastic rattle for toddlers. The rattle is generally marketed through exclusive retailers located in upscale shopping malls. The company used 23 retailers. In late 2016, Diana Suares, the president of the company, was considering alternative marketing plans presented to her by Bill Dings, the marketing manager. Based on sales from January through October 2016, Diana expected that 2016sales would amount to 300,000 units at $8.00 per unit. Diana Suares also had with her some cost data for 2016 supplied by the CFO, Don Capp. Don expects that these costs are reliable estimates for a volume up to 400,000 units. Beyond 400,000 units, the company would have to rent additional machines (with a capacity of 100,000 units each) at an annual cost of $50,000 per machine (in 2016). The data are presented in Exhibit A. Total fixed manufacturing and selling and administrative costs are each expected to increase in 2017 by 10 percent because of inflation. Variable costs per unit and the selling price would stay the same as in 2016. Exhibit A (2016 Cost Data) Manufacturing costs for rattles (based on production volume of 300,000 units).

Direct material $0.80 per unit *

Direct labor $10 per hour * (Each worker can make 20 units in an hour)

Packaging $0.75 per unit *

Power, supplies, indirect labor, and other variable production costs $1.20 per unit *

Supervisory salaries, equipment rental, and miscellaneous production costs $1.80 per unit *

These costs vary in total with production volume

Selling and administrative costs (based on sales volume of 300,000 units).

Commissions to sales staff 10% of price to retailer **

Shipping $0.50 per unit **

Advertising and promotion $0.60 per unit

Administrative staff salaries, depreciationon office equipment, etc $0.90 per unit **

These costs vary in total with sales volume.

1. What is the materials cost at 300,000 units?

2. What is the materials cost at 400,000 units?

3. What is the total variable cost at 200,000 units?

4. What is the total variable cost at 400,000 units?

5. What is the fixed manufacturing cost at 300,000 units?

6. What is the fixed manufacturing cost at 400,000 units?

7. What is the fixed manufacturing cost at 450,000 units?

8. What is the total fixed cost at 300,000 units?

9. What is the total fixed cost at 400,000 units?

10. What is the contribution margin per unit?

11. What is the contribution margin if 1000 units are sold?

12. What is the contribution margin if100,000 units are sold?

13. What is the profit at the break even volume?

14. What is the contribution margin at the break even volume?

15. What is the break even volume?

16. What is the profit at 1000 units above break even?

17. What is the profit at 2000 units below the break even?

18. What is the profit at 100,000 units above break even?

19. What is the margin of safety at 1000 units above break even?

20. What is the sales volume required to give $100,000 profit before tax?

21. What is the sales volume required to give $180,000 after tax, assuming a tax rate of 40%?
Answer #1

1. Materials cost at 300,000 units = 300000 x $0.80 = $240,000

2. Materials cost at 400,000 units = 400000 x $0.80 = $320,000

3. Variable Cost at 200000 units

Direct Material $        160,000 =200000*0.8
Direct Labor $        100,000 =200000*10/20
Packaging $        150,000 =200000*0.75
Power, supplies, indirect labor, and other variable production costs $        240,000 =200000*1.2
Commissions to sales $        160,000 =200000*8*10%
Shipping $        100,000 =200000*0.5
Total Variable Costs $        910,000

4. Variable Cost for 400000 units

Direct Material $        320,000 =400000*0.8
Direct Labor $        200,000 =400000*10/20
Packaging $        300,000 =400000*0.75
Power, supplies, indirect labor, and other variable production costs $        480,000 =400000*1.2
Commissions to sales $        320,000 =400000*8*10%
Shipping $        200,000 =400000*0.5
Total Variable Costs $     1,820,000

5. Fixed manufacturing cost at 300,000 units

Supervisory salaries, equipment rental, and miscellaneous production costs $        594,000 =300000*1.8*1.1
Total Fixed Manufacturing Costs $        594,000


6. Fixed manufacturing cost for 400,000 units

Supervisory salaries, equipment rental, and miscellaneous production costs $        594,000 =300000*1.8*1.1
Total Fixed Manufacturing Costs $        594,000


7. Fixed manufacturing cost for 450,000 units

Supervisory salaries, equipment rental, and miscellaneous production costs $        594,000 =300000*1.8*1.1
Rent for additional machines $           50,000
Total Fixed Manufacturing Costs $        644,000

8. Total Fixed cost for 300,000 units

Supervisory salaries, equipment rental, and miscellaneous production costs $        594,000 =300000*1.8*1.1
Advertising and promotion $        198,000 =300000*0.6*1.1
Administrative staff salaries, depreciation on office equipment $        297,000 =300000*0.9*1.1
Total Fixed Costs $     1,089,000

9. Total Fixed Cost for 400,000 units

Supervisory salaries, equipment rental, and miscellaneous production costs $        594,000 =300000*1.8*1.1
Advertising and promotion $        198,000 =300000*0.6*1.1
Administrative staff salaries, depreciation on office equipment $        297,000 =300000*0.9*1.1
Total Fixed Costs $     1,089,000

As per HOMEWORKLIB RULES, we are supposed to solve 4 parts, i have done 9, kindly post other parts separately